3. (15 points) You are a monopolist with the following cost and demand conditions: P = 100 – 2Q and C(Q) = 50 + Q2.

a. (5 points) Determine the profit-maximizing output and price.

b. (5 points) Graph this solution and show your profits and the deadweight loss to society in your graph.

c. (5 points) Determine the actual amount of deadweight loss.
Answer #1

Monopoly: @ P=100 - 20 TR= p.a - 100 - 200 TR = 100Q २२२ MR = CT2 da = 100 CV - na =foo -4Q sota² MC = dTC d +20 २२ Monopolis

4Q+ R Q = 100 62 a 6 Q 16.67 unit. Substituting for into Qm into p equation p = 100.-2C16.67) =log -33.33 = 66.66 P = $66.66